OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = −3(1/3)^x + 1 − 2?

OpenStudy (anonymous):

@mathstudent55 @mathmate

OpenStudy (anonymous):

@satellite73

OpenStudy (mathmate):

Is the question \(f(x)=-3(\frac{1}{3})^x +1-2 \) or \(f(x)=-3(\frac{1}{3})^x +x-2 \)